OpenStudy (jordanloveangel):

A line has a slope of 7 and contains the point (5, −9). Which equations represent the line? Choose all answers that are correct. A. −7x + y = −44 B. 7x – y = 44 C. (y − 9) = 7(x + 5) D. (y + 9) = 7(x − 5)

OpenStudy (jordanloveangel):

@iGreen

OpenStudy (horsegirl27):

I can help

OpenStudy (jordanloveangel):

@horsegirl27

OpenStudy (jordanloveangel):

please do

OpenStudy (horsegirl27):

First of all I think it's B or C let me work on it. You try it too and tell me what u get

OpenStudy (jordanloveangel):

k

OpenStudy (igreen):

We use the point-slope form.. \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value of the point, \(x_1\) is the x-value of the point, and \(m\) is the slope. Here, \(y_1\) is -9, \(x_1\) is 5, and \(m\) is 7. Plug them in what do you get?

OpenStudy (jordanloveangel):

i got B

OpenStudy (igreen):

@horsegirl27 Please don't give answers..we have to guide the asker to the answer.

OpenStudy (jordanloveangel):

but idk the second answer

OpenStudy (horsegirl27):

ok sorry u already explained so I was confirming the final answer ;)

OpenStudy (igreen):

B is incorrect..

OpenStudy (jordanloveangel):

but using the point slope form you'll have: (y-y1) = m(x-x1) y - (-9) = 7(x -5) y + 9 = 7(x - 5) or y + 9 = 7x - 35 7x - y = 9 + 35 7x - y = 44 or multiplying the whole equation with -1 7x - y = 44 -7x + y = -44

OpenStudy (jordanloveangel):

???????????????????

OpenStudy (igreen):

Yep, -7x + y = -44 is correct.. And so is y + 9 = 7(x - 5)

OpenStudy (igreen):

y + 9 = 7(x - 5) Distribute 7 into the parenthesis: y + 9 = 7x - 35 Subtract 9 to both sides: y = 7x - 44 Subtract 7x to both sides: -7x + y = -44

OpenStudy (jordanloveangel):

so A and D?

OpenStudy (jordanloveangel):

@undeadknight26

OpenStudy (jordanloveangel):

@fallenangelorchid

OpenStudy (jordanloveangel):

@sleepyjess

OpenStudy (fallenangelorchid):

using the point slope form you'll have: (y-y1) = m(x-x1) y - (-9) = 7(x -5) y + 9 = 7(x - 5) or y + 9 = 7x - 35 7x - y = 9 + 35 7x - y = 44

OpenStudy (fallenangelorchid):

or multiplying the whole equation with -1 7x - y = 44 -7x + y = -44

OpenStudy (jordanloveangel):

k thx

OpenStudy (jordanloveangel):

but theres 2 right answers

OpenStudy (fallenangelorchid):

Yeah, i'd go with A and D.

OpenStudy (jordanloveangel):

k thx soooo much